Mastery Problem: Activity-Based Costing

WoolCorp

WoolCorp buys sheep’s wool from farmers. The company began operations in January of this year, and is making decisions on product offerings, pricing, and vendors. The company is also examining its method of assigning overhead to products. You’ve just been hired as a production manager at WoolCorp.

Currently WoolCorp makes two products: (1) raw, clean wool to be used as stuffing or insulation and (2) wool yarn for use in the textile industry.

The company would like you to evaluate its costing methods for its raw wool and wool yarn.

Single Plantwide Rate

WoolCorp is currently using the single plantwide factory overhead rate method, which uses a predetermined overhead rate based on an estimated allocation base such as direct labor hours or machine hours. The rate is computed as follows:

Single Plantwide Factory Overhead Rate = (Total Budgeted Factory Overhead) ÷ (Total Budgeted Plantwide Allocation Base)

WoolCorp has been using combing machine hours as its allocation base.

The company would like to consider activity-based costing. In order to understand their current system better, you evaluate WoolCorp’s current method of costing for raw wool and wool yarn. The production staff has compiled the following information for you on the production of 500 pounds of either raw wool or wool yarn:


Factory
Overhead Type
Budgeted
Factory
Overhead
Sorting $25,600   
Cleaning 38,400   
Combing 1,400   
Raw Wool Wool Yarn
Hours of combing machine use required 70 30

In the following table, use combing machine hours as the allocation base for assigning overhead costs to each product. When required, round your answers to the nearest dollar.

Single Plantwide Factory Overhead Rate: $ per combing hour

Raw Wool Wool Yarn
Allocated factory overhead cost $ $

Activity-Based Costing

In order to compare WoolCorp’s current method with activity-based costing, you interview the production staff and compile the following information, which relates to the costs for raw wool and wool yarn.

Type of Cost Activity Base Total Cost
Sorting Hours of sorting $25,600
Cleaning Units of cleaning machine power 38,400
Combing Hours of combing machine use 1,400
Raw Wool Wool Yarn
Hours of sorting required 1,000     4,000    
Units of cleaning machine power required 1,800     4,200    
Hours of combing machine use required 70     30    

In the following table, compute and enter the activity rate for each of the three activities. If required, round your answers to the nearest cent.

Activity Activity Rate
Sorting $ per sorting hour
Cleaning $ per unit of cleaning machine power
Combing $ per hour of combing machine use

In the following table, allocate the costs of sorting, cleaning, and combing based on the rates of activity consumed by each product’s process. When required, round your answers to the nearest dollar.

Raw Wool Wool Yarn
Sorting cost $ $
Cleaning cost
Combing cost
Total cost $ $
Answer #1

Single Plantwide Factory Overhead Rate: $65400/100 = 654 per combing hour

Raw Wool Wool Yarn
Allocated factory overhead cost $654*70 = 45780 $19620

In the following table, compute and enter the activity rate for each of the three activities. If required, round your answers to the nearest cent.

Activity Activity Rate
Sorting $25600/5000 = 5.12 per sorting hour
Cleaning $38400/6000 = 6.40 per unit of cleaning machine power
Combing $1400/100 = 14 per hour of combing machine use

In the following table, allocate the costs of sorting, cleaning, and combing based on the rates of activity consumed by each product’s process. When required, round your answers to the nearest dollar.

Raw Wool Wool Yarn
Sorting cost $5120 $20480
Cleaning cost 11520 26880
Combing cost 980 420
Total cost $17620 $47780